OpenStudy (anonymous):

Passing through (-3,1) 3x-2y=5. write answer in point slope and slope intercept form?

OpenStudy (anonymous):

are you looking for a line parallel to 3x-2y=5? cause I actually didn't understand the question. SORRY.

OpenStudy (anonymous):

yes parallel

OpenStudy (anonymous):

OK. Slope intercept form is basically in the form y=mx+b. So rewrite the equation 3x-2y=5, which will become y=(3/2)x-(5/2). Point slope form is in the form y-y1=m(x-x1). So substitute the points (-3, 1) for x1 and y1, where the slope m=3/2, which you'll get y-1=3/2(x+3), which is the point slope form. Hope this helps!

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

i had y-1=3/2(x+3)

OpenStudy (anonymous):

so that's where i went wrong

OpenStudy (anonymous):

Yup, that's the point slope form. You're right!

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

so, it was y = 3/2x - 5/2 to make a line parallel to it, you will start with y = 3/2x because parallel lines have the same slopes. then you will find a y intercept that will satisfy the equation when you substitute (-3,1). then the equation would be 1 = -4.5 + ?. so you will find a number that will make it 1 = 1. 5.5 is the suitable number. the answer would be y = 3/2x + 5.5 im pretty sure this is right. i used a graphing program to help with and they end up parallel to 3x-2y=5.